if a vortex of a triangle E(1,1) & the middle point of the sides through it (-2,3)&(5,2) find the other vorties

Grade:Upto college level

2 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

Dear Satya Padhy,

 

the ratio middle points divides two vertex is 1:1

 

using formula      (x1*1 +x2*1)/(1+1) = -2       here   E(1,1)=(x1,y1)  and say A(x2,y2) B(x3,y3)

 

substitute x1 here   we get   x2=-4-x1 = -4-1 =-5

 

y2= 3-y1 = 3-1 =2

 

(x2,y2)=(-5,2)

 

x3= 5-x1= 5-1 =4

y3=2-1= 1

 

(x3,y3)=(4,1)

 

 


Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay

 

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Satya,

 

 given middle point so it divides the line into 1:1

so,  -2 = (1 + x1)/2 or x1 = -5

and,  3 = (1 + y1)/2  or y1 = 5

so one of the vertex is (-5,5)

similarly find the other one.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay